OpenStudy (anonymous):

can someone help me with this question: A ladder is leaning against a wall. It is 5 ½ feet from the wall and has an angle of elevation 55°. How long is the ladder?

OpenStudy (anonymous):

5'5?

OpenStudy (campbell_st):

draw a diagram and show the information  now decide with trig ratio will allow you to find the length of the ladder

OpenStudy (anonymous):

sin is O/H cos is A/H and tan is O/A

OpenStudy (campbell_st):

so the ladder will be the hypotenuse.... so do you have the opposite side or the adjacent side...?

OpenStudy (anonymous):

adjacent

OpenStudy (anonymous):

65 inches

OpenStudy (campbell_st):

ok... so you have \[\cos(55) =\frac{65}{H}\] now make H the subject so you can find the length of the ladder

OpenStudy (anonymous):

yeah Im working it out now thanks i think i got it

OpenStudy (anonymous):

where are we getting the 65? I think we are wrong

OpenStudy (anonymous):

65 inches or 6'5inches?

OpenStudy (campbell_st):

12 inches = 1 ft so 5 ft = 5* 12 = 60 and so its 60 + 5 = 65 inches...